### v3

# XpookyNet: Advancement in Quantum System Analysis through Convolutional Neural Networks for Detection of Entanglement

Abstract
The applications of machine learning models in quantum information theory have witnessed a surge in recent years, driven by the sophistication of quantum computing (QC). However, relying on inappropriate models has often resulted in suboptimal accuracy. A custom deep convolutional neural network (CNN), XpookyNet, has been developed to address this. XpookyNet is explicitly tailored for quantum systems analysis and aims to detect fundamental quantum attributes, such as entanglement in two- and three-qubit systems and potentially more qubit systems. It harnesses the power of CNNs to provide instantaneous and meticulous results. With an accuracy of 98.53% achieved in merely a few epochs, XpookyNet effectively handles the inherent complexity of quantum information, providing deeper insights into QC. The study also investigates quantum features and their relation to the purity of a density matrix directly related to noise in NISQ-era quantum computation. Preparing the density matrix in a format compatible with customary CNNs is vital in dissecting quantum systems. Quantum-applied CNNs must also be refined utilizing recent methods and extensive metrics to enhance and ensure their performance.
preprint: APS/123-QED
I Introduction
In quantum mechanics, an uncanny phenomenon known as quantum entanglement arises when two or more particles interact so that their quantum states become related in a way that the neither state $\psi_{AB}\neq\psi_{A}\otimes\psi_{b}$ nor $\langle A\otimes B\rangle\neq\langle A\rangle\otimes\langle B\rangle$ [1, 2]. Instantaneous changes in one particle are simultaneously mirrored in others, irrespective of their spatial separation, with the non-locality condition of $P(A=a,B=b)\neq P(A=a)\cdot P(B=b)$ holding [3]. Quantum Computing needs entanglement to form and rise, as it allows the system to perform multiple calculations simultaneously using many qubits, achieves secure and high key rates, and underpins quantum annealing and variational quantum algorithms to solve intractable problems and optimization problems [4, 5, 6, 7, 8].
The first step in manipulating systems is to detect the presence and amount of entanglement, so various entanglement detection criteria have been proposed [9, 10, 11]. However, the positive partial transpose (PPT) criterion can determine entanglement with certainty for systems $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ where the Hilbert space is $\mathcal{H}_{AB}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ with the necessary size of $2\otimes 2$ or $2\otimes 3$ non-mixed states. In other words, Hilbert space includes entangled mixed states that still meet the PPT conditions called bound-entangled states [12].
Comprehensively assessing quantum entanglement requires choosing the right criterion. While criteria like concurrence, negativity, and relative entropy of entanglement have advantages, the entanglement of formation (EoF) stands out for its operational interpretation of the resources needed to create a specific state. Unlike its counterparts, it provides a deeper understanding of entanglement phenomenon [13, 14, 15].
Quantifying and classifying entanglement beyond three qubits is an NP-Hard problem. The generalized concurrence still measures entanglement, but its calculation is computationally intensive for mixed states and it cannot pinpoint individual qubits [16]. Entanglement witnesses are another practical tool to detect entanglement in many-body quantum systems. A witness $W$ is a Hermitian operator that for all separable states, $\langle W\rangle_{\rho_{sep}}=\text{Tr}(W\rho_{sep})\geq 0$ but for some entangled states $\langle W\rangle_{\rho_{ent}}=\text{Tr}(W\rho_{ent})<0$ [17]. It detects entanglement without fully characterizing the system or performing a tomography; however, due to the exponential growth of variables with higher qubit counts, it requires optimization in high-dimensional spaces. Quantum witnesses can be optimized using machine learning (ML) because they can quickly identify patterns in large datasets, making them ideal for solving complex problems [18, 19]. Illustrations in Fig. 1 give a grasp of compartmentalizing methods for classifying states.
Figure 1: Entangled or separable in Hilbert space. (a) Adjustment of a witness as a linear hyperplane and its optimization failure. (b) Entanglement witness optimization approaches, including linear: from $W_{1}$ to $W_{1}^{\prime}$, and nonlinear: from $W_{2}$ to $W_{2}^{\prime}$ [20, 21]. (c) Hindrance of precision due to the convexity of the target space. (d) Improvement of naive ML usage as convex witnesses isolator resulting in an ungeneralized model.
Among ML techniques, CNNs are versatile and robust, transforming research fields and impacting our daily lives [22]. They can attain accurate results on any dataset if it is properly encoded and the model is adapted to the data [23].
A link between ML and quantum information (QI) has been extensively studied recently [24, 25, 26, 27, 28]. There has yet to be a model with rigorous accuracy for two-qubit entanglement detection problem, and the design of CNN models for QI applications remains relatively unexplored. The research needs to be more comprehensive in characterizing quantum system states beyond two qubits under the entanglement types. Additionally, the impact of noise on state identification regarding mixed states has yet to be extensively explored.
Fig. 2 shows the research outline of this study. Initially, we create random density matrices and use them for training CNN. The density matrices produced for the dataset are categorized using the entanglement of formation (EoF). EoF underlies classifying the matrices based on their characteristics. The matrices are then pre-processed into an extended-tensor format to prepare them for subsequent processing, as explained in section II. In section III, we develop custom CNN models. The CNN model is designed and refined through a training and optimization process. Plateau machination, bracing layers, and branching features are added to enhance the model’s structure. The model is then subjected to testing to evaluate its performance. The resulting classification outcomes are used to assess the model’s applicability and make necessary adjustments. This approach ensures an accurately designed, trained, optimized, and tested CNN model. Section IV involves constructing confusion a matrix to assess the balance and symmetry of the results. It helps understand the model’s accuracy and errors made. Key metrics like ACC, MCC, and FNR are calculated to extensively analyze the study’s performance and reliability.
Figure 2: A comprehensive overview of the study’s scheme.
II Quantum Data
There are two methods to combine quantum and classical computers for ML. The first method involves using a quantum computer to process and transform quantum data into a form that classical ML algorithms can understand [29]. The second method transfers data, like density matrix, from the quantum to the classical domain [30], which is being addressed in subsection . Both methods are in the early stages of research but have the potential to offer significant benefits.
II.1 Density Matrix and Entanglement
Multi-party quantum systems exhibit quantum entanglement between qubits. The density matrix is a Hermitian matrix that allows us to expand its states and probabilities. Generally, it is defined as
$\rho=\sum\nolimits_{i=1}^{N}\lambda_{i}|\Psi_{i}\rangle\langle\Psi_{i}|,$ (1)
where, $\sum_{i}\lambda_{i}=1$ For all states $|\Psi\rangle$. It represents a multi-party quantum system, especially when it is in a mixed state. So, they allow us to calculate properties such as entanglement and coherence of an actual quantum system and obtain their expectation values and time evolution [31, 32]. Hence, the equation for the expectation value in pure states given as $\langle\hat{A}\rangle=\langle\psi|\hat{A}|\psi\rangle$, turns into $\langle\hat{A}\rangle=\mathrm{Tr}(\hat{\rho}\hat{A})$. Similarly, the evolution equation for pure states described as $\frac{d}{dt}|\psi(t)\rangle=\frac{1}{i\hbar}\hat{H}(t)|\psi(t)\rangle$, is transformed into $\frac{d}{dt}\hat{\rho}(t)=\frac{1}{i\hbar}[\hat{H}(t),\hat{\rho}]$, where the relative density matrix of $|\psi\rangle$ is represented by $\hat{\rho}$.
In this study, density matrices are used as data to enable CNNs to extract features from quantum systems. We labeled the data based on the EoF, which for $\rho$ of a system is calculated:
$EoF=\inf{\left[\sum_{i}p_{i}S(\rho_{i})\right]},$ (2)
where the infimum is taken over all possible state decompositions into a probabilistic mixture of pure product states. The von Neumann entropy of the reduced density matrix for subsystem A is calculated by:
$S(\rho_{A})=-\mathrm{Tr}\left[\rho_{A}\log_{2}{(\rho_{A})}\right],$ (3)
where $\rho_{A}$ represents the reduced density matrix of subsystem A.
II.2 Constructing Extended-Tensor
QI relies on complex numbers, but ordinary CNNs cannot accommodate them. We convert the density matrix into a three-dimensional tensor. Let $M$ be the initial $n\times n$ matrix constituted of complex numbers, and $M_{ij}=a_{ij}+\text{i}b_{ij}$ is defined as the complex number in the $i_{\text{th}}$ row and $j_{\text{th}}$ column of $M$. As illustrated in Fig. 3, we separate the density matrices into two matrices: one containing real numbers
$\sum_{j=0}^{n}\sum_{i=0}^{n}A_{ij}=\Re(M_{ij})$ (4)
and the other containing imaginary numbers
$\sum_{j=0}^{n}\sum_{i=0}^{n}B_{ij}=-\Im(M_{ij})*\text{i}$ (5)
Finally, we obtain a tensor that forms the input data as follows:
$\sum_{k=0}^{1}\begin{cases}C_{ijk}=A_{ij},&k=0\\ C_{ijk}=B_{ij},&k=1\end{cases}$ (6)
Tensor $C$ congregates matrix A and B, which have separated the real and imaginary parts into different layers of the 3D on each row $i$ and column $j$.
Figure 3: The process of converting complex number data to real number data.
II.3 Two-qubit data generation
To generate the density matrix for a two-qubit system, we generate a random complex matrix $M$ of size $4\times 4$ with elements of $M_{i,j}=a+ib$ where $a$ and $b$ are drawn from a Gaussian distribution. A Hermitian matrix $H$ is formed from the random complex matrix $M$ by calculating $H=M+M^{\dagger}$ to ensure that $H$ is Hermitian. Then $\text{Tr}(H^{2})>0$ is checked to ensure $\langle\psi|\rho|\psi\rangle\geq 0$. Then the density matrix $\rho$ is formed from the Hermitian matrix $H$ by normalizing it, i.e., $\rho=H/\text{Tr}(H)$, to ensure that the $\text{Tr}(\rho)=1$. QuTiP facilitates this procedure [33]. We regulate our density matrices to have 75% non-zero elements so it is more likely to generate a complex, mixed quantum state and feature significant quantum superposition. However, since begetting random two-entangled $\rho$ is approximately three times more likely than begetting separable $\rho$, it is vital to balance our dataset to prevent bias and improve CNN’s accuracy.
We generate one million $4\times 4$ random density matrices as our dataset [34]. This dataset contains 500,000 entangled and 500,000 separable $\rho$ that are explained in Appendix.A. The matrices are labeled by calculating the EoF of each matrix using the Qiskit [35].
$\text{Label}(\rho)=\begin{cases}\text{Ent},&\text{EoF}(\rho)>0\\ \text{Sep},&\text{EoF}(\rho)=0\end{cases}$ (7)
The amount of entanglement is also stored thanks to the EoF function.
II.4 Three-qubit data generation
Figure 4: An overview of three-qubit: (a) An allegory of dividing state space into categories. (b) Quantum circuits yielding distinct entangled systems: (b.1) Fully entangled state $\rho_{ABC}$, (b.2) Partial entangled states $\rho_{AB|C}$, (b.3) Partial entangled states $\rho_{BC|A}$, (b.4) Partial entangled states $\rho_{AC|B}$.
In the three-qubit quantum system Hilbert space onwards, it is not pervasive to only divide states into entangled or separate. As depicted in Fig. 4(a), entangled states itself are divided into four types $|\Psi_{ABC}\rangle$, $|\Psi_{AB|C}\rangle$, $|\Psi_{BC|A}\rangle$, and $|\Psi_{AC|B}\rangle$. We detailed in Appendix.B how to obtain each type but superficially saying, partial entanglements $|\Psi_{AB|C}\rangle$, and $|\Psi_{BC|A}\rangle$ are provided with a $\otimes$ operator between an entangled $|\Psi_{i,j}\rangle$ and $|\Psi_{k}\rangle$ but $|\Psi_{AC|B}\rangle$ is not that easy. Also, $|\Psi_{ABC}\rangle$ are demanding, which puts our focus on known three-qubit states GHZ, W state, and Graph state. Fig. 4(b) physically demonstrates the entangled states’ type, which exerts randomness by single random unitary operations $U_{r}^{A,B,C}$ and entanglement by $U^{Q}_{Ent.}$ operator, where $Q$ determines monolith entangled qubits. The number of $Q$ over $n$ qubits is as follows:
$q=\sum_{i=2}^{n}\binom{n}{i},$ (8)
when $n=3$ qubit, $q$ is 4. $Q$ includes {$AB|C$, $BC|A$, $AC|B$, $ABC$}.
We create 250,000 density matrices for three-qubit states, divided into five balanced categories as a dataset, which takes only 102 seconds to prepare 100% pure states when running on a regular CPU.
III Designing a CNN
CNNs consist of convolutional layers with multiple kernels that extract patterns from low to high-level features, allowing complicated patterns within a matrix to be extracted.
In CNN, the 3D convolution operation is expressed as:
$Y_{i,j,k}=\sigma\left(b+\sum_{p=0}^{P-1}\sum_{q=0}^{Q-1}\sum_{r=0}^{R-1}X_{i+p,j+q,k+r}\times W_{p,q,r}\right),$ (9)
where $Y_{i,j,k}$ is the output element at position $(i,j,k)$, $X_{i+p,j+q,k+r}$ is the input element at position $(i+p,j+q,k+r)$, $W_{p,q,r}$ is the weight parameter at position $\left(p,q,r\right)$, $b$ is the bias term, and $\sigma$ is the activation function. In equation (9), $P$, $Q$, and $R$ are the kernel dimensions that work as a filter. The sum over $p$, $q$, and $r$ represents the convolution operation, where the kernel slides over the input tensor and filters the input tensor’s corresponding elements. The activation function introduces non-linearity into the layers, allowing the network to learn complex patterns in high-dimensional space.
Figure 5: XpookyNet’s whole scheme. Blue layers show the principal artery and green layers show the branches. The last layer consists of two dense layers.
Entanglement detection requires a high-capacity CNN with many layers. However, longer sequences of layers increase the chance of vanishing gradients of back-propagation. This is caused by activation functions that map input values to small intervals. Leaky ReLU, defined as:
$f(x)=\begin{cases}ax,&x<0\\ x,&\text{else}\end{cases},$ (10)
is a solution to vanishing gradients as it introduces a slight negative slope $a$ for values below zero, allowing for continued learning.
To understand CNN performance, a loss function is needed. The loss function categorical cross entropy (CCE) is calculated for a given number of classes, denoted as $C$.
$L_{CCE}=-\sum_{i=1}^{C}{y_{i}\log{(\hat{y_{i}})}}$ (11)
In CCE’s Equation (11), $y_{i}$ represents the actual label for class $i$, and $\hat{y_{i}}$ denotes the predicted SoftMax probability for class $i$. Considering only entangled or not, the proposed equation (2) simplifies to a binary classification problem by setting $C=2$. The modified equation becomes:
$-\sum_{i=1}^{2}y_{i}\log{(\hat{y_{i}})}=-y_{1}\log{(\hat{y_{1}})}-y_{2}\log{(\hat{y_{2}})}$ (12)
Equation (12) is further simplified to represent the formula for binary cross entropy as follows:
$L_{BCE}=-y\log{(\hat{y})}-(1-y)\log{(1-\hat{y})}$ (13)
The back-propagation algorithm optimizes weight and bias in equation (9) over loss functions:
$\delta_{j}^{k}\equiv\frac{\partial a_{j}^{k}}{\partial E}$ (14)
Here, $\delta_{j}^{k}$ is the loss at layer $k$ for node $j$, $a_{j}^{k}$ is the activation of node $j$ at layer $k$, and $E$ is the loss function. The algorithm is used to update the weights of the network during training. The loss $\delta_{j}^{k}$ at layer $k$ is dependent on the loss $\delta_{k}^{k+1}$ at the next layer $k+1$. Thus, loss flows backward from the last layer to the first layer.
A CNN is designed and developed by incorporating and assembling these components [36].
III.1 Proposed CNN
CNN requires layers to be appropriately arranged and select hyper-parameters wisely, such as the number of kernels and filters in each layer. We set different kernel sizes, as shown in Fig. 5. Kernel size selection depends on the area it covers in the input tensor. Smaller kernels, e.g., $2\times 2$ detect detailed patterns and larger kernels, e.g., $4\times 4$ detect broader patterns [37].
The proposed CNN, XpookyNet, can determine the spooky action at a distance, its amount, and the entangled qubits by using varied activation functions in the last layer accordingly.
XpookyNet has parallel layers with expanded kernel sizes, and it uses conventional convolutional layers that can be replaced with more efficient ”separable convolutional layers.” Batch Normalization (BN) stabilizes CNNs and improves their performance, accuracy, and speed [38]. Separable convolutions make CNN training easier, faster, and more accurate while reducing computational complexity and overfitting. However, they may be less effective with low-dimension tensors and require more layers to achieve accuracy, similar to traditional CNNs [39].
III.2 Training CNNs
XpookyNet uses the stochastic gradient descent (SGD) optimizer with momentum and saves the most accurate model with minimum loss. The learning rate is adjusted when the loss reaches a plateau [40]. Parallel paths help prevent vanishing gradients again by giving a shortcut to backward flow [41]. Also, balancing and shuffling the data improves results.
When training models, it is crucial to select the learning rate carefully. The learning rate determines how much the model changes in response to the error estimate. It is recommended to begin with a high learning rate as it enables the model to learn faster. The learning rate is reduced manually when the loss function stops improving, allowing the model to make more miniature adjustments. This process is repeated throughout the training to guide the model towards better performance [42, 43].
IV Results Evaluation
Figure 6: The proposed XpookyNet performance in the learning process. (a) Accuracy, and (b) Loss function graphs for the methods listed
Table 1: Overview of Methods used in this study for two-qubit entanglement detection. Conv., Brch., Sep., and Plat. stand for convolutional, branched model, separable convolution, and ReduceLROnPlateau, respectively.
Model Structure
Methods
ACC
FNR
MCC
Epoch Time
Number of Conv. Layer
Number of Parameters
NN
0.8325
0.1826
0.6603
19s44ms
0
3,233
Simple Conv.
0.9500
0.0622
0.9189
51s55ms
10
1,334,913
Brch.
0.9624
0.377
0.9276
63s22ms
14
2,015,521
BN. Sep.
0.9632
0.0504
0.9205
57s20ms
10
1,080,263
Plat.
0.9789
0.0244
0.9576
48s33ms
10
1,334,913
Brch. Plat. 0.9852 0.0156 0.9712 64s58ms 14 2,015,521
Plat. BN. Sep.
0.9749
0.0440
0.9452
60s47ms
10
1,080,263
Brch. BN. Sep.
0.9664
0.0518
0.9315
75s62ms
14
1,324,615
Brch. Plat. BN. Sep.
0.9761
0.0440
0.9452
75s95ms
14
1,335,430
We assess XpookyNet for detecting entanglement in quantum systems. Our analysis includes loss function and accuracy graphs, overview tables, and classification performance using relevant metrics:
Accuracy (ACC) measures the ratio of correctly predicted instances to the total number of instances in the dataset.
$\small ACC=\frac{TP+TN}{TP+TN+FP+FN}$ (15)
where True Positives ($TP$) is the number of correctly predicted positive instances, True Negatives ($TN$) is the number of correctly predicted negative instances, False Positives ($FP$) is the number of incorrectly predicted positive but are negative, False Negatives ($FN$) is the number of incorrectly predicted negative but are positive.
False Negative Rate (FNR) measures the proportion of actual positive instances that were incorrectly predicted as negative.
$\small FNR=\frac{FN}{TP+FN}$ (16)
Matthews Correlation Coefficient (MCC) is a balanced measure that takes into account true and false positives and negatives.
$\small MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (17)
IV.1 Two-qubit Entanglement Detection
Results of 10,000 random two-qubit test states are presented in Table 1 for entanglement detection in two qubits, which is crucial for building and characterizing larger quantum systems and their states [44].
The NN is a baseline with the lowest ACC, MCC, and a notable FNR. Simple Convolution improves ACC and MCC but retains a discernible FNR. The Branched CNN enhances accuracy but results in a higher FNR than Simple Convolution, albeit with an improved MCC. BN and Separable convolution layers achieve a commendable balance with high ACC, a lower FNR, and an impressive MCC. The Reducing the learning rate on plateaus (ReduceLROnPlateau) outperforms all others with the highest ACC, MCC, and minimal FNR. Combining branched CNN with the ReduceLROnPlateau method is the most optimal choice.
In only 14 epochs, XpookyNet (”Brch.Plat.”) achieves a 98.53% accuracy. The network can even determine the entanglement degree with mean absolute error 0.0034. Graphs 6 (a) depict the performance evolution of methods over epochs. The NN starts with high accuracy but flats around 0.8 ACC. In comparison, convolutional models show progressive improvement, especially the ones that exploit the ReduceLROnPlateau technique. In Graph 6 (a), the combination of the branched model with the ”ReduceLROnPlateau” configuration (”Brch.Plat.”) stands out for its significant improvement over time. Graphs 6 (b) illustrate the loss function of the methods to showcase the effect of ReduceLROnPlateau on epochs $8$ and $12$, emphasizing its action on the loss function’s plateau.
Strategic model selection and combination are crucial for optimal performance. Combining models can lead to continued improvement, emphasizing the need for a nuanced approach to model architecture and configuration.
Figure 7: Based on purity values of 100%, 83%, 56%, and 37%, confusion matrices illustrate the classification performance of the CNN model in three-qubit quantum systems.
IV.2 Three-qubit Entanglement Detection
XpookyNet utilizes separable convolution layers and BNs in a three-qubit setting for enhanced performance on larger tensors compared to traditional convolutional layers.
Improving XpookyNet becomes essential, especially for analyzing quantum states at different purity levels to address noise, errors, and imperfections in the NISQ era.
Table 2: Overview of metrics used to assess three-qubit entanglement detection.
Classes
Purity
ACC
FNR
MCC
Sep.
1
0.9980
0.0000
0.9938
0.83
0.9951
0.0035
0.9848
0.56
0.9731
0.0275
0.9197
0.37
0.9191
0.0605
0.7616
$AB|C$
1
0.9994
0.0030
0.9981
0.83
0.9986
0.0060
0.9956
0.56
0.9888
0.0405
0.9648
0.37
0.9652
0.1395
0.8889
$A|BC$
1
0.9994
0.0030
0.9981
0.83
0.9987
0.0050
0.9959
0.56
0.9932
0.0244
0.9784
0.37
0.9722
0.1090
0.9117
$AC|B$
1
0.9991
0.0035
0.9972
0.83
0.9983
0.0076
0.9946
0.56
0.9900
0.0400
0.9685
0.37
0.9718
0.1110
0.9104
$ABC$
1
0.9997
0.0015
0.9991
0.83
0.9987
0.0045
0.9959
0.56
0.9938
0.0205
0.9806
0.37
0.9795
0.0605
0.9356
Overall 1 0.9978 0.0022 0.9972
0.83
0.9947
0.0053
0.9934
0.56
0.9694
0.0306
0.9618
0.37
0.9039
0.0961
0.8799
XpookyNet (”Brch. Plat. BN. Sep.”) classifies three-qubit systems with high accuracy on density matrices with purity 1, but its performance declines as purity decreases, depicted in Fig. 7. Owing to imperfections and noise, real-world quantum systems lose their entanglement as purity drops. Thus, the purity plunge primarily affects XpookyNet’s classification of partial entangled states that encompass characteristics of separable and entangled states, with an average 11.7% performance shrinkage. As purity decreases, fully entangled states can appear partial entangled or separable, leading to misclassification in XpookyNet. Mixing density matrices with random ones can also blur class boundaries. However, XpookyNet performs strongly, with only a 5.98% performance decrease on average, despite purity drop effects in these two classes. This behavior of XpookyNet aligns with the physical intuition about quantum entanglement and the effects of system imperfections, demonstrating the network’s ability to capture the complex dynamics of quantum systems despite noise and imperfections [45].
XpoolyNet’s performance declines across all classes as purity decreases from 1 to 0.37. Table 2 expresses ACC and MCC decrease while FNR increases, indicating misclassifications. For separable states, purity decline results in a 7.9% ACC decrease and a 23.4% MCC drop. The FNR increases by 0.0605, a notable rise in errors, especially FP. TN and TP remain relatively accurate, pointing to entanglement features obscuring results. Partial entangled states ($AB|C$, $A|BC$, $AC|B$) show less drastic ACC and MCC decrease than separable states. However, FNR increases by an average of 62.5%, indicating significant errors at decimated purity. Fully Entangled (ABC) states experienced a negligible impact on purity decrease, with only a 1.83% ACC drop and a 6.36% MCC decrease. Oddly, three-qubit full entanglement detection functions better than two-qubit problem due to the limited known tripartite entangled set of GHZ, W, and graph states.
V Conclusion and Discussion
We propose XpookyNet, a novel approach using CNNs to detect entanglement in quantum systems. Our model represents density matrices as 3D tensors and employs parallel layers with expanded kernel sizes for improved efficiency. Throughout our experimentation, we systematically evaluate different model architectures and configurations. The results demonstrate that XpookyNet outperforms its counterparts regarding evaluation metrics, effectively detecting entanglement quantum systems. Our analysis also extends to the impact of purity levels on XpookyNet’s prowess. The model’s mastery subsides as purity drops, although it performs flawlessly on high-purity matrices. Nevertheless, in tandem with classifying the quantum system’s conditions, the growth in classes’ errors shows alignment with physical intuition about quantum entanglement, confirming XpookyNet’s robustness and connoisseur of quantum idiosyncrasies. By entrenching a procedure that generates all Hilbert space’s partial entangled states $AB|C$, $AC|B$, and $BC|A$ and a deft model that captures the quantum system’s traits through its density matrix, the detection of tripartite entanglement ABC becomes a cinch since recognizing that the density matrix of a fully entangled $ABC$ system encapsulates the integrals derived from the partial entangled states $AB|C$, $AC|B$, and $BC|A$. By analogy, more qubit entanglement configurations are paved[46, 47]. Moreover, XpookyNet can be considered a criterion for quantum systems due to its high accuracy and understanding of quantum features.
Acknowledgements.
We extend our heartfelt appreciation to the University of Tehran for fostering a collaborative and nurturing environment that greatly benefited the researchers engaged in this study.
Appendix A Two-qubit density matrix
Two-qubit separable mixed states are expressed:
$\rho_{sep}=\sum_{i=1}^{m}{\lambda_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}},$ (18)
where $\sum_{i}\lambda_{i}=1$ and $0\leq\lambda_{i}\leq 1$, summing over $i=1$ to $m$. $\rho_{i}^{A}$ and $\rho_{i}^{B}$ are arbitrary density matrices of A and B qubits respectively. Two-qubit entangled mixed states are expressed as:
$\rho_{ent}=\sum_{i=1}^{m}{\lambda_{i}\rho_{i}^{AB}},$ (19)
$\rho_{i}^{AB}$ is the arbitrary density matrix of the two-qubit entangled state.
Appendix B Generation of Three-qubit density matrix
In the three qubits case, two-qubit entangled states are required to build partial entangled states.
B.1 Three-qubit Separable density matrix
Separable mixed states are expressed:
$\rho_{sep}=\sum_{i=1}^{m}{\lambda_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}\otimes\rho_{i}^{C}},$ (20)
where $\sum_{i}\lambda_{i}=1$ and $0\leq\lambda_{i}\leq 1$, with $m$ iterating from 1 to an arbitrary number. The greater the value of $m$, the less pure the state.
B.2 Partial entangled density matrix
Partial entangled mixed states of a three-qubit state are produced in three cases.
Case 1: Partial entangled pairs A and B are generated by:
$\rho_{AB|C}=\sum_{i=1}^{m}{\lambda_{i}\rho_{i}^{AB}\otimes\rho_{i}^{C}},$ (21)
Where $\sum_{i}\lambda_{i}=1$ and $0\leq\lambda_{i}\leq 1$, with $m$ iterating from 1 to an arbitrary number. $\rho^{AB}$ is a density matrix of a two-qubit entangled state.
Case 2: Partial entangled pairs B and C are generated by:
$\rho_{A|BC}=\sum_{i=1}^{m}{\lambda_{i}\rho_{i}^{A}\otimes\rho_{i}^{BC}},$ (22)
Where $\sum_{i}\lambda_{i}=1$ and $0\leq\lambda_{i}\leq 1$, $\rho_{i}^{BC}$ is an arbitrary density matrix of a two-qubit entangled state.
Case 3: Partial entangled pairs A and C are generated by state vectors of the entangled state $|\psi_{AC}\rangle$ and separated state $|\psi_{B}\rangle$ considered as:
$|\psi_{AC}\rangle=[a_{0}\ \ a_{1}\ \ a_{2}\ \ a_{3}]^{T}$ (23)
and
$|\psi_{B}\rangle=[b_{0}\ \ b_{1}]^{T}.$ (24)
Therefore, the partial entangled states are considered as:
$\begin{split}|\psi_{AC|B}\rangle=[a_{0}b_{0}\ \ a_{1}b_{0}\ \ a_{0}b_{1}\ \ a_{1}b_{1}\\ a_{2}b_{0}\ \ a_{3}b_{0}\ \ a_{2}b_{1}\ \ a_{3}b_{1}]^{T},\end{split}$ (25)
where $a$ and $b$ are corespondent magnitudes of equations (23) and (24). Ultimately, the partial entangled states are generated by:
$\rho_{AC|B}=\sum_{i=1}^{m}{\lambda_{i}(|\psi_{B|AC}\rangle_{i}\langle\psi_{B|AC}|_{i})},$ (26)
where $\sum_{i}\lambda_{i}=1$ and $0\leq\lambda_{i}\leq 1$ and $|\psi_{B|AC}\rangle$ is the generated partial entangled state.
B.3 Three-qubit GHZ state
The state vector of a random three-qubit state that GHZ is applied to is described as:
$|\Psi_{GHZ}\rangle=\frac{1}{\sqrt{N_{GHZ}}}\left(\cos{(\epsilon)}|000\rangle+\sin{(\epsilon)}e^{i\phi}|\Phi_{ABC}\rangle\right)$ (27)
with initial states:
$|\Phi_{ABC}\rangle=|\varphi_{A}\rangle|\varphi_{B}\rangle|\varphi_{C}\rangle$ (28)
$|\varphi_{A}\rangle=\cos{(\theta_{A})}|0\rangle+e^{i\phi_{A}}\sin{(\theta_{A})}|1\rangle,$ (29)
$|\varphi_{B}\rangle=\cos{(\theta_{B})}|0\rangle+e^{i\phi_{B}}\sin{(\theta_{B})}|1\rangle,$ (30)
$|\varphi_{C}\rangle=\cos{(\theta_{C})}|0\rangle+e^{i\phi_{C}}\sin{(\theta_{C})}|1\rangle,$ (31)
where
$N_{GHZ}=1/(1+\cos{(\delta)}\sin{(\delta)}\cos{(\alpha)}\cos{(\beta)}\cos{(\phi)}).$
[…]
B.4 Three-qubit W-state
A random state vector that undergoes the W-state operation can be written as:
$\left|\Psi_{\text{W}}\right\rangle=$ $\frac{1}{\sqrt{N_{\text{W}}}}(a\ |001\rangle+b\ |010\rangle+c|100\rangle-d\ |\phi\rangle),$
where
$N_{W}=1/\sqrt{|a|^{2}+|b|^{2}+|c|^{2}+|d|^{2}},$
is the normalization factor and $|\phi\rangle$ is a superposition of remaining states with W-state.
B.5 Three-qubit Graph state
The graph state can be considered as:
$\left|\Psi_{\text{Graph}}\right\rangle=$ $\frac{1}{\sqrt{N_{\text{Graph}}}}(\alpha_{0}\ |000\rangle+\alpha_{1}\ |001\rangle$
$\quad+\alpha_{2}\ |010\rangle-\alpha_{3}\ |011\rangle+\alpha_{4}\ |100\rangle$
$\quad+\alpha_{5}\ |101\rangle-\alpha_{6}\ |110\rangle+\alpha_{7}\ |111\rangle),$ (33)
where normalization factor is
$N_{Graph}=1/\sqrt{|\alpha_{0}|^{2}+\ldots+|\alpha_{7}|^{2}}.$